OpenStudy (anonymous):

Silas took 15 bags of glass to the recycling center. He still has 7 bags of plastic to take to the recycling center. Which equation could be used to find x, the total number of bags of glass and plastic Silas will take to the recycling center? A. x + 15 = 7 B. x - 15 = 7 C. 15 - x = 7 D. x - 7 = -15

OpenStudy (texaschic101):

I am gonna go with B. x (original amount) --- took 15 bags x - 15 still has 7 x - 15 = 7 x = 15 + 7 x = 22